Circular No. 6135

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995D IN NGC 2962 S. Benetti, C. Mendes de Oliveira, European Southern Observatory

(ESO); and A. Manchado, Instituto de Astrofisica de Canarias, report that inspection of a fully-reduced CCD spectrogram (range 350–100 nm, resolution 1 nm) obtained on Feb. 13.18 UT with ESO 1.5-m telescope (+ Boller & Chivens spectrograph), confirms that SN 1995D is a type-Ia supernova, ∼ 1 week before maximum light. The spectrum is dominated by P-Cyg lines of intermediate-mass elements superimposed on a very blue continuum. Visible lines include Ca II (H and K; infrared triplet); Si II (635.5 nm), S II (546.8, 561.2, 565.4 nm) and faint O I (777.3 nm). Strong lines of Si III (456.0, 574.0 nm) are also present, as in spectra of similar phase of SN 1990N. Si II (597.2 nm) is unusual weak, if present at all. The expansion velocities deduced from the minimum of Si II 635.5-nm and Ca II H and K lines, corrected by the redshift (0.00655) of the parent galaxy, are ∼ 10900 and 18200 km/s, respectively.

NOVA AQUILAE 1995 T. Iijima, H. Esenoglu, and L. Rosino, Asiago Astrophysical Observa-

tory, report: “Optical spectrograms of N Aql 1995 were obtained on Feb. 10.2 UT with the Asiago 1.2-m telescope (+ prismatic spectrograph; range 410–510 nm, resolution 0.2 nm/pixel at 450 nm). The spectra show prominent emissions lines of H I, Fe II (multiplet 24,28,32,37,38,42), and some weak lines of [Fe II] and [Fe III]. The FWHMs of H I lines are ∼ 1300 km/s, while those of Fe II are 1400 km/s, and none has a P-Cyg profile. These features indicate that this nova is in the early decline stage.”

Visual magnitude estimates: Feb. 12.436 UT, 9.0 (J. Bortle, Stormville, NY); 13.445, 9.4 (Bortle); 14.18, 9.0 (L. Kiss, Szeged, Hungary).

SUPERNOVA 1994Y IN NGC 5371 L. Wang, Z.-W. Li, A. Clocchiatti, and J. C. Wheeler, McDonald Ob-

servatory, communicate: “A high-resolution (0.07 nm/pixel) spectrogram of SN 1994Y taken at the McDonald 2.1-m telescope on Feb. 8 shows a narrow absorption feature (full width ∼ 365 km/s) at 662.2 nm, superimposed on top of the strong Hα line. This signifies a significant amount of circumstellar material. Observations at all wavelengths are encouraged.”

1995 February 15 Daniel W. E. Green